### v1

# Triple Phase Transitions: Understanding the Learning Dynamics of Large Language Models from a Neuroscience Perspective

Abstract
Large language models (LLMs) often exhibit abrupt emergent behavior, whereby new abilities arise at certain points during their training. This phenomenon, commonly referred to as a “phase transition”, remains poorly understood. In this study, we conduct an integrative analysis of such phase transitions by examining three interconnected perspectives: the similarity between LLMs and the human brain, the internal states of LLMs, and downstream task performance. We propose a novel interpretation for the learning dynamics of LLMs that vary in both training data and architecture, revealing that three phase transitions commonly emerge across these models during training: (1) alignment with the entire brain surges as LLMs begin adhering to task instructions (Brain Alignment and Instruction Following), (2) unexpectedly, LLMs diverge from the brain during a period in which downstream task accuracy temporarily stagnates (Brain Detachment and Stagnation), and (3) alignment with the brain reoccurs as LLMs become capable of solving the downstream tasks (Brain Realignment and Consolidation). These findings illuminate the underlying mechanisms of phase transitions in LLMs, while opening new avenues for interdisciplinary research bridging AI and neuroscience.
1 Introduction
Large language models (LLMs) often exhibit abrupt emergent behaviors, whereby new abilities arise from scaling the model size, the amount of training data, or the number of training steps (Wei et al., 2022a). This behavior has predominantly been discovered through evaluations of model outputs (e.g., downstream task performance), leading to numerous breakthroughs (Wei et al., 2022b; Caballero et al., 2023; Du et al., 2024). Recent efforts in mechanistic interpretability have begun to uncover the underlying internal changes that occur during training process, aiming to elucidate these emergent phenomena (Olsson et al., 2022; Chen et al., 2023). However, prior studies have largely examined these transitions in isolated perspectives, making it unclear how these transitions, along with other additional aspects, interact with each other.
In seeking to clarify the underlying principles of deep neural networks (DNNs), neuroscience research has yielded human-centered insights into DNNs through comparisons between their internal representations and human brain activity (Yamins et al., 2014; Güçlü & van Gerven, 2015). This approach has recently been extended to LLMs (Jain & Huth, 2018; Goldstein et al., 2022; Oota et al., 2022), revealing that brain activity has a stronger alignment with larger- and higher-performance models (Antonello et al., 2023). These findings highlight the potential of the human brain, an intricate system with diverse functions, to serve as a tool for model evaluation. Because most existing studies have only examined trained models, however, they offer limited insight into how emergent phenomena arise during the learning process.
In this study, we focus on the learning dynamics of LLMs in an attempt to provide a comprehensive understanding of emergent phenomena by integrating three analytical perspectives: brain encoding analysis, which assesses alignment with human brain activity; probing analysis, which detects shifts in internal representations; and benchmark analysis, which measures downstream task performance [^1] All experimental resources (source code, training data, and model checkpoints) will be publicly available to promote reproducibility and future research. See Sections A.1 and A.3 for the details.. We demonstrate that multiple LLMs, each characterized by distinct architectures and training data, exhibit a robust, common three-stage phase transition in their learning dynamics, including the precise timing of these transitions. Specifically, (1) alignment with the entire brain surges once the LLMs begin to follow downstream task instructions (Brain Alignment and Instruction Following); (2) surprisingly, the LLMs diverge from the brain during a period in which their downstream task accuracy temporarily stagnates (Brain Detachment and Stagnation); and (3) alignment with the brain reoccurs once LLMs become capable of solving the downstream tasks (Brain Realignment and Consolidation). Although prior work has reported stronger brain alignment with larger and higher performance models (Antonello et al., 2023), our findings reveal that the learning trajectory of LLMs is more complex than previously thought, consisting of multiple phases. By examining multiple LLMs trained on distinct language datasets, we highlight the influence of the training data on these learning dynamics. Taken together, our findings demonstrate that incorporating human brain activity as a biologically grounded benchmark reveals how the emergent capabilities of LLMs form and consolidate during training, offering essential insights for safer, more interpretable language models.
Figure 1: Overview of the study. a Brain encoding analysis. b Probing analysis (top) and the benchmark analysis (bottom). c Three phase-transition phenomena during the learning process of LLMs, as identified through the results of the encoding, probing, and benchmark analyses. Red, green, and blue lines indicate encoding, probing, and benchmark accuracies, respectively, accros all LLM checkpoints.
2 Related Work
Phase Transitions in LLMs
LLMs acquire distinct abilities throughout their training process as their model size becomes larger, the amount of training data rises, and the number of training steps increases. While some abilities emerge gradually (Kaplan et al., 2020; Hoffmann et al., 2022), others appear abruptly (Ganguli et al., 2022; Wei et al., 2022a). Previous studies have identified “emergent abilities”, such as the onset of chain-of-thought prompting beyond a certain scale (Wei et al., 2022a). To understand LLMs internally, mechanistic interpretability seeks to unveil the internal computations learned by neural networks (Elhage et al., 2021; Dai et al., 2022). Internal phase transitions in the learning process have also recently been observed in this field (Olsson et al., 2022; Chen et al., 2023; Park et al., 2024), shedding light on how abrupt performance gains align with shifts in internal representations. For example, Olsson et al. (2022) found “phase changes” in Transformers, where induction heads suddenly arise and enable extended contexts to be handled through in-context learning. However, prior studies have largely focused on transitions in isolated perspectives, leaving it unclear how these shifts, alongside other relevant aspects, interact.
Interpretability from Neuroscience Insights
From a neuroscience perspective, related work has compared the internal representations of DNNs with human brain activity (Yamins et al., 2014; Güçlü & van Gerven, 2015), demonstrating that the hierarchical structure of high-performing DNNs mirrors the hierarchical processing of the human visual cortex. More recent efforts have employed linear mappings from LLM internal representations to brain activity, probing language processing similarities between LLMs and the human brain (Jain & Huth, 2018; Schrimpf et al., 2021; Goldstein et al., 2022; Oota et al., 2022). Notable examples include comparisons of learning characteristics between LLMs and the human brain (Millet et al., 2022; Caucheteux & King, 2022; Caucheteux et al., 2023; Aw et al., 2023; Aw & Toneva, 2023; Antonello et al., 2023; Tuckute et al., 2024a; b; Antonello & Huth, 2024; de Varda et al., 2025) and discussions of the changes in latent dimensionality within LLMs (Cheng & Antonello, 2024).
3 Methods
3.1 Large Language Models
We analyze the learning dynamics of LLMs from three distinct perspectives. For this, we use four pre-trained models with available training checkpoints: OLMo-2 (OLMo et al., 2024), OLMo-0724 (Groeneveld et al., 2024), and LLM-jp (LLM-jp et al., 2024) for the main analysis, and Amber (Liu et al., 2023) for the control analysis. Table 1 presents an overview of these LLMs. OLMo-2, OLMo-0724, and Amber are English-centric LLMs trained on publicly accessible datasets, while LLM-jp is a bilingual Japanese-English model trained on a roughly equal mix of Japanese and English. Each model uses a different training corpus and tokenizer, with vocabulary sizes of 100,352 (OLMo-2), 50,304 (OLMo-0724), 99,584 (LLM-jp), and 32,000 (Amber). These LLMs have between 6.74–7.3B parameters, 32 hidden layers, and a hidden dimension of 4,096. They are all based on a decoder-only Transformer architecture (Vaswani et al., 2017), but each LLM incorporates a few critical modifications. There are also notable differences across these LLMs in terms of their layer normalization and attention mechanisms (see Section A.1 for architectural details). The multilayer perceptron (MLP) layers require most parameters and represent essential features (Bereska & Gavves, 2024; Geva et al., 2021). Thus, we use their neural activation as the internal representations of each model. We use 28 checkpoints for OLMo-2 (1B-3,896B training tokens), 23 for OLMo-0724 (4B–2,724B), 27 for LLM-jp (4.2M–1,258B), and 18 for Amber (3.5B–1,259B). See Section A.1 for further details on the model sources and specific checkpoints.
3.2 Brain Encoding Models
Our initial approach for analyzing the learning dynamics of the LLMs involves an investigation of how their activations progressively align with brain activity during the training process. Specifically, for each checkpoint described in Section 3.1, we perform an encoding analysis by evaluating the prediction accuracy of a learned linear mapping from each layer’s activations to brain activity (Naselaris et al., 2011; Nishimoto et al., 2011; Huth et al., 2012) (Figure 1a). These analysis are conducted separately for each participant.
fMRI Datasets
We use the Narrative Movie fMRI Dataset (Yamaguchi et al., 2024; Nakagi et al., 2024) for our brain encoding analysis. This dataset provides brain activity data from six healthy participants with normal or corrected-normal vision (three females; ages 22–40, mean = 28.7), while they freely watched 8.3 hours of movies or drama series inside a 3T functional Magnetic Resonance Imaging (fMRI) scanner. All participants were right-handed and native Japanese speakers. The dataset includes nine video clips of movies or drama series as stimuli: eight international titles and one Japanese title. The international titles were dubbed into Japanese, allowing the participants to view all clips in Japanese. The dataset also contains three types of natural language annotation from the videos. We use the Narrative Content annotation, which describes the background story of the scene at 5-s intervals, for the main analysis and the Objective Information annotation, which describes the objects in the scene every second, for the control analysis, both in English and Japanese. Section A.2 provides additional details regarding the dataset. In this study, we use 29,993 seconds of data covering all six participants. We divide the data into training and test datasets. All prediction performance results are computed using the test dataset. Specifically, we use the fMRI scanning sessions corresponding to the last split of each movie or drama series (7,737 seconds in total) as the test data. The remainder of the sessions (22,262 seconds in total) constitutes the training data.
Model Construction
We extract the activations of the LLMs for the annotations from each hidden layer. They consist of several tokens for each time point. We then average the activations across tokens. Because multiple annotations exist for each second, we extract the activations for each annotator and then average them across all annotators. We then train an L2-regularized linear model to predict the voxel-level brain activity from the corresponding activations of the LLM neurons. We estimate the model weights from the training data, then apply them to the test data. The regularization parameters are determined for each voxel by cross-validation during training. We then evaluate the prediction accuracy by computing the Pearson’s correlation coefficients between the predicted and measured fMRI signals. Statistical significance is assessed using a blockwise permutation test that compares the correlations between predicted and measured signals against the correlations obtained after shuffling the measured signals. We set the threshold for statistical significance to $p<0.05$ and correct for multiple comparisons using the FDR procedure. We model the hemodynamic delay in the BOLD signal, assumed to be 8–10 seconds. See Section A.3 for additional details about model construction and region-of-interest (ROI) analyses. In the main analysis, we focus on later layers within each LLM. Specifically, we focus on layer 25 in both OLMo-2 and LLM-jp, and layer 30 in OLMo-0724. These layers show the most substantial checkpoint-wise changes and are strongly involved in the emergent behavior of each model, as verified in all three evaluations (encoding, probing, and benchmark analyses; see Section A.6 for details). We confirm that the layers surrounding the main focal layer exhibit similar patterns of results (see Figure B.5).
3.3 Probing
As the second approach, we investigate how downstream task-relevant representations are progressively acquired within the LLMs during the training process. To this end, we perform a probing analysis by evaluating how well linear models can predict neuron activations from the answer labels of downstream tasks across all hidden layers (Figure 1b, top). These analysis are conducted separately for each LLM checkpoint.
Downstream Datasets
We use four downstream tasks: Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2020), CommonsenseQA (CSQA) (Talmor et al., 2019), AI2 Reasoning Challenge (ARC) (Clark et al., 2018), and HellaSwag (Zellers et al., 2019). MMLU assesses broad knowledge and problem-solving abilities using multidisciplinary coverage of 57 subjects, CSQA tests everyday conceptual commonsense reasoning, ARC probes elementary-level scientific knowledge, and HellaSwag assesses contextual commonsense reasoning in typical scenarios. For both the probing and benchmark analyses (see Section 3.4), we use 5-shot prompts in both English and Japanese. We exclude prompts exceeding each model’s maximum context length. See Section A.4 for additional details about downstream datasets. In the probing analysis, each dataset is split into training and test datasets at a 4:1 ratio. Because MMLU comprises multiple subject areas, we split the dataset by subject. All prediction performance results are computed using the test dataset. Examples of the prompts used in these datasets are presented in Section A.7.
Probing for MLP Activations
We first feed the downstream task prompts into the LLMs and extract the final-token neuron activations in each layer. For each sample in the downstream tasks, we create an answer matrix (the number of samples $\times$ the number of choices), where the correct and incorrect choices are labeled as 1 and 0, respectively. Our objective is to learn a mapping from this answer matrix to the neuron activations across all layers of the LLMs. To this end, we use L2-regularized linear regression to estimate the linear weights that transform the answer matrix corresponding to the training data into the observed neuron activations in the LLMs, and then apply these learned weights to the test data. The regularization parameter is optimized via 4-fold cross-validation for each neuron in the training dataset. To mitigate the subject-based bias of MMLU, we shuffle the training indices, and then perform cross-validation to ensure balanced distributions in each fold. We then evaluate the prediction accuracy by computing the Pearson’s correlation coefficients between the predicted and actual activations. The amount of information necessary for each task that is retained at each LLM checkpoint is quantitatively evaluated at the neuron level. In this study, we focus on the same layers as in the encoding analysis (Section 3.2).
3.4 Benchmark
As the third approach for analyzing the learning dynamics of LLMs, we investigate how the ability to solve downstream tasks emerges as the training process continues. Specifically, we use each LLM checkpoint to evaluate performance on the downstream task datasets from Section 3.3 (Figure 1b, bottom). The evaluation metric is the ratio of correctly answered items, i.e., the fraction of samples for which the model outputs exactly match the correct answers. We use the llm-jp-eval library [^2] https://github.com/llm-jp/llm-jp-eval for this analysis.
3.5 Direct Analysis of MLP Activations
To investigate how internal representations in LLMs change over the course of training, we compare their properties across checkpoints using the correlation and dimensionality of activations. We compute the Pearson’s correlation coefficients of the activations across LLM checkpoints. Specifically, we use the same dataset as for the encoding analysis, extract the activations from the same layers, and calculate their correlations across different checkpoints. We also quantify the dimensionality of the activations at each checkpoint and examine changes across checkpoints. Here, we adopt the Intrinsic Dimension (ID), which has garnered attention as a means of characterizing the nature of LLM internal representations. We estimate the IDs at each checkpoint by applying the Generalized Ratios Intrinsic Dimension Estimator (GRIDE) (Denti et al., 2022) to each layer’s activations of the same dataset used for the encoding analysis (see Section A.5 for details).
4 Results
Figure 2: Learning dynamics of LLMs exhibiting three phase transitions. The horizontal axis denotes the number of training tokens. The vertical axis denotes the average encoding accuracy for all voxels of a single participant (DM06) (red lines), the benchmark accuracy (blue lines), and the average probing accuracy for all LLM neurons calculated using MMLU (green lines). We select layers 25, 30, and 25 from OLMo-2, OLMo-0724, and LLM-jp, respectively, to capture the transitions that occur at each phase of the learning dynamics. The background color indicates the LLM phase. The legend indicates whether the language has been learned sufficiently by the model. No checkpoints preceding the 109 training tokens have been made publicly available aside from LLM-jp.
4.1 Learning Dynamics of LLMs
[…]
Figure 2 presents the results for OLMo-2, OLMo-0724, and LLM-jp when using MMLU for the probing and benchmark analyses. The brain encoding result shows the average prediction accuracy across all brain voxels inferred from LLM neurons in the specified layer, whereas the probing result shows the average prediction accuracy across all LLM neurons. Section B.1 provides analogous results for adjacent layers, other downstream tasks (CSQA, ARC, HellaSwag), other LLMs (Amber), other languages (Chinese), and other annotations.
Figure 3: Changes in the relationship with the brain. Differences in encoding accuracy among checkpoints for three participants (DM06, DM03, and DM07), projected onto the inflated (top, lateral, and medial views) and flattened cortical surface (occipital areas are at the center, only for DM06), for both the left and right hemispheres. We have chosen OLMo-2’s checkpoints that capture the transitions at each phase of the learning dynamics. Brain regions with significant accuracy at either of the two checkpoints are colored ($p<0.05$, FDR-corrected). Voxels that exhibit higher accuracy at the later checkpoint are indicated in red, whereas those that exhibit higher accuracy at the earlier checkpoint are indicated in blue.
4.2 What Happens During Each Phase?
Section 4.1 characterized the three phase transitions that arise during LLM training, along with interpretive insights drawn from encoding, probing, and benchmark analyses. Nevertheless, the exact internal changes that occur during each training phase remain unclear. To gain a deeper understanding and provide further interpretation, we examine how the relationship with the brain changes at the voxel level and how the internal representations in the LLMs shift at the neuron level in each phase.
Changes in the Relationship with the Brain
To characterize the voxel-level relationship between LLMs and the brain across the three phases, we calculate the difference in brain encoding accuracy for each voxel between two representative training checkpoints that capture each phase transition, and project these changes onto each participant’s cerebral cortex. Figure 3 shows that, using the activations of OLMo-2, the voxel-level accuracy patterns shift distinctly from one phase to another. Specifically, in Phase 1—where the average prediction accuracy increases sharply—the accuracy gains are broadly distributed throughout the cerebral cortex, especially the temporal, occipital, and frontal cortex. By contrast, in Phase 2—marked by a decrease in overall brain encoding accuracy—the changes tend to vary more on a voxel-by-voxel basis, with some voxels within the temporal cortex still showing improvement. In Phase 3, the accuracy rises again across broad brain regions (see Figure B.11). Furthermore, comparing the post-transition prediction accuracies in Phase 3 with those in Phase 1 reveals that, in certain voxels within the temporal cortex, the post-Phase 3 accuracies are higher than the post–Phase 1 accuracies (see Figure B.11).
Overall, these findings indicate that once the LLMs begin following the task instructions (Phase 1), their internal representations exhibit better global alignment with brain activity. During the subsequent stagnation phase (Phase 2), this global alignment partially recedes, and after Phase 3, the LLMs realign with the brain, reinforcing the similarity to brain regions involved in semantic processing, in particular. Notably, these patterns only emerge when the input language matches one on which the model has been trained. Further details for all participants, ROIs, other phases, and LLMs (OLMo-2, OLMo-0724, LLM-jp) can be found in Figures B.11B.16.
Figure 4: Probing results. a Evolution of probing accuracy for English/Japanese MMLU throughout the training process, assessed at layers 5, 15, 25, and 30 of OLMo-2. The horizontal axis indicates the probing accuracy. The vertical axis indicates the number of neurons that fall within each 0.01 accuracy bin. The legends corresponds to the number of training tokens. b Relationship between probing accuracy in the neurons of OLMo-2 across English MMLU, CSQA, and ARC. Each axis denotes the probing accuracy for the respective task, and the color gradient reflects the number of training tokens. The legend shows the correlation coefficient between certain two tasks.
Evolution of LLM Internal Representations for Downstream Tasks
We now investigate how the neuron-level internal representations evolve over the course of training. For this purpose, we examine changes in the distribution of probing accuracy for each neuron at the early, middle, and final checkpoints (105B, 1259B, 3896B training tokens in OLMo-2) of Phase 3. Figure 4a presents the results for neurons in layers 5, 15, 25, and 30 of OLMo-2 when using MMLU. From these results, we find that, in the English MMLU setting, later-layer neurons progressively increase in probing accuracy, indicating that, as the LLMs acquire downstream task proficiency, they also develop more specialized neuron activations. Figures B.17B.19 show the results corresponding to all downstream tasks for OLMo-2, OLMo-0724, and LLM-jp, confirming that a similar trend emerges across every downstream task when each model is provided with the language on which it has been trained.
Figure 5: The Nature of Activations. a Variations in correlation coefficients of the activations of OLMo-2 across checkpoints. b IDs (purple line) and average encoding accuracy for all voxels of a single participant (DM06) (red line) across checkpoints.
We can also explore whether different sets of neurons acquire specialized representations for distinct downstream tasks, focusing on the same checkpoints examined in Figure 4a. Figure 4b illustrates the relationships among the per-neuron probing accuracies of OLMo-2 for the English MMLU, CSQA, and ARC. Across all three tasks, it is evident that certain neurons begin to manifest representations specialized for each downstream task as training progresses. Intriguingly, some neurons develop such specialized representations for all tasks, while others remain specific to only a subset of tasks. Moreover, the per-neuron accuracies for MMLU and ARC are highly correlated (r = 0.864), whereas those for CSQA demonstrate weak correlations with the other two tasks (r = 0.029 with MMLU, r = 0.027 with ARC). Furthermore, focusing on OLMo-0724—which, much like OLMo-2, demonstrates high benchmark accuracy on downstream task performance—we confirm that OLMo-0724 yields results that are analogous to those of OLMo-2 (Figure B.20). This implies that substantial transformations, specific to each task’s characteristics, are emerging within these models. The results for all downstream tasks, including HellaSwag, are discussed in Section B.3.
4.3 Changes in the Nature of Activations
Based on brain encoding and probing analyses, the preceding results have demonstrated that shifts in neuronal activation occur over the course of training. We now examine whether these alterations reflect fundamental shifts in the activations of the LLMs themselves.
As shown in Figure 5a, within-phase activations are highly similar, but substantial changes emerge at each phase transition. Interestingly, although the brain encoding accuracy does not vary dramatically between Phases 2 and 3, the underlying activations differ considerably. Next, we quantify the dimensionality of these activations using the IDs defined by GRIDE. Figure 5b shows that there is a strong similarity between the brain encoding accuracy and IDs during Phases 1 and 2, consistent with earlier findings that single-checkpoint encoding accuracy and dimensionality are highly correlated (Cheng & Antonello, 2024). Our findings extend this result by demonstrating a similarly robust correlation across multiple checkpoints. Nevertheless, the brain encoding accuracy and dimensionality are not perfectly congruent (see Phase 3), leaving open questions about the aspects of activation changes that intrinsic dimensionality alone does not capture.
5 Discussion and Conclusions
In this study, we interpreted the learning dynamics of LLMs from three perspectives: alignment between the brain and the LLMs, internal representations associated with downstream tasks, and actual downstream task performance. Our novel approach elucidated a three-stage phase transition—including the precise timing of each transition—that emerges consistently during training, even among LLMs that diverge substantially in both training data and architecture.
The Brain Alignment and Instruction Following Phase and Brain Realignment and Consolidation Phase resemble the emergent phenomena reported by Wei et al. (2022a; b); Caballero et al. (2023); Olsson et al. (2022), whereby models abruptly gain novel abilities after passing a threshold. Prior work has largely focused on changes in outputs or the emergence of specific mechanisms. By integrating human brain alignment and shifts in LLM internal representations, we have shown that these factors progress in unison, highlighting multi-scale indicators behind abrupt gains and stagnations in model performance. Furthermore, the marked emergence of these phenomena in languages extensively represented within the model’s training corpus implies that deep internalization of their statistical patterns can trigger significant shifts in both brain alignment and overall performance. Hence, the presence or absence of these phase transitions may indicate whether an LLM has adequately assimilated and comprehended a language’s statistical structure.
Additionally, the learning dynamics that we have revealed are reminiscent of the “synaptic overproduction and pruning” posited in neuroscience, where an initial synaptic overgrowth eventually refines the neural circuits (Peter R., 1979; Bourgeois et al., 1994). These transient changes in the brain suggest the potential for more efficient neural information processing, leading to enchanced cognitive function. Thus, a promising avenue for future work involves examining whether the phenomenon observed in this study arises from the LLM pruning superfluous internal representations in pursuit of more refined ones.
This study has also revealed that voxel-level brain alignment does not simply increase over the training process, but fluctuates across distinct phases. Previous research linking model representations to brain activity (Yamins et al., 2014; Güçlü & van Gerven, 2015; Schrimpf et al., 2021; Caucheteux et al., 2023) has not fully addressed such learning dynamics. While Antonello et al. (2023) reported a rise in alignment with increasing model size, our findings depict a more intricate path: an initial surge in brain alignment, a decline, and a final resurgence. This suggests that LLMs may adopt distinct computational strategies at various stages, rather than gradually acquiring brain-like language representations. Moreover, alignment with semantic regions in the temporal cortex generally rises over the course of training, implying a dynamic reorganization in how models correspond to brain activity.
By capturing shifts in the activations of LLM neurons pertaining to downstream tasks, we elucidated how task-specific neurons form and how much specialization they share among tasks. These findings align with previous studies showing neurons that encode particular linguistic information (Tenney et al., 2019; Dai et al., 2022; Wang et al., 2022; Gurnee et al., 2023). We observed that neurons contributing substantially to certain tasks emerge abruptly in later layers as training progresses. Furthermore, the variations in neuron sharing across tasks suggest that differences in activation patterns—reflecting task-specific characteristics (e.g., required capabilities, difficulty levels, and answer formats)—are indeed discernible at the neuronal level. Taken together, our results offer a more comprehensive view of LLM neurons during the learning process, building on past work that focused solely on trained models. Further in-depth analyses of these neurons would be a fascinating avenue for future research.
Finally, we extended the findings of Cheng & Antonello (2024), who demonstrated a positive correlation between the dimensionality of LLM activations in trained models and the alignment of those LLMs with the brain. Our results confirm that this correlation remains robust during training, suggesting that the observed phase transitions are tied to changes in activation dimensionality. It is crucial to note that although these metrics confirm that fundamental transformations occur in the representation space, they do not fully explain the specific underlying mechanisms. By examining how these transformations relate to both the brain and downstream tasks, as in this study, a more comprehensive and nuanced understanding of the underlying processes can be achieved.
In summary, this study has elucidated the existence of three distinct phase transitions during LLM training, as evidenced by three perspectives: alignment with the human brain, shifts in internal representations pertinent to downstream tasks, and task performance. We interpreted these transitions as indispensable internal transformations that enable the model to acquire downstream task capabilities. This study is the first to show that changes in brain alignment, internal representation, and model performance advance in tandem by leveraging the human brain—the only known system (aside from LLMs) capable of processing complex language. These findings highlight the critical importance of examining multiple signals to gain a comprehensive understanding of LLM learning dynamics, including emergent phenomena. Furthermore, they suggest the potential to harness human brain activity in the pursuit of explainable and human-aligned language models.
Acknowledgements
Y.T. was supported by PRESTO Grant Number JP-MJPR23I6. S.N. was supported by KAKENHI JP24H00619 and JST JPMJCR24U2.
Author Contributions
All authors conceived the study. Y.N., K.T., S.Y., and Y.T. analyzed the data. Y.N. wrote the original draft. Y.N. and Y.T. wrote the manuscript with the consultation by the other authors. All authors reviewed the manuscript.
Appendix A Additional Methods
A.1 Large Language Models
We used the allenai/OLMo-2-1124-7B, allenai/OLMo-7B-0724-hf, llm-jp/llm-jp-3-7.2b, and LLM360/Amber models available on Hugging Face for OLMo-2, OLMo-0724, LLM-jp, and Amber. All checkpoints of LLM-jp will be made publicly available, although it has only released its final checkpoint. Tables 1 and 2 present an overview of the LLMs and detailed information on their respective training checkpoints used in this study. In selecting these checkpoints, we took particular care to ensure that the number of training tokens was as closely aligned as possible across the four LLMs.
Model Layers Width Params. Trn. Tokens (Ckpts.)
OLMo-2 32 4096 7.3B 1B - 3896B (28)
OLMo-0724 32 4096 6.89B 4B - 2724B (23)
LLM-jp 32 4096 7.29B 4.2M - 1258B (27)
Amber 32 4096 6.74B 3.5B - 1259B (18)
Table 1: Overview of LLMs
Model Checkpoints
OLMo-2 150, 600, 900, 2K, 3K, 4K, 5K, 10K, 15K, 20K, 25K, 35K, 45K, 55K, 65K, 72K, 150K, 230K, 300K, 370K, 441K, 513K, 584K, 656K, 727K, 799K, 870K, 928.646K (Training step)
OLMo-0724 1K, 2K, 2.5K, 3.5K, 4.5K, 5K, 10K, 15K, 20.5K, 25.5K, 35.5K, 45.5K, 55.5K, 65K, 72K, 149.5K, 230K, 300K, 370K, 442K, 514K, 585K, 649.65K (Training step)
LLM-jp 1, 2, 4, 8, 20, 30, 60, 100, 300, 500, 1K, 2K, 3K, 4K, 5K, 10K, 15K, 20K, 25K, 35K, 45K, 55K, 65K, 72K, 150K, 230K, 300K (Training step)
Amber 1, 2, 3, 4, 5, 6, 12, 18, 24, 30, 42, 54, 66, 78, 86, 179, 275, main (Checkpoint)
Table 2: Details of the training checkpoints
The architectural variations among the LLMs used in this study encompass layer normalization, activation functions, positional embeddings, and attention mechanisms. All models are derived from a decoder-only Transformer (Vaswani et al., 2017) architecture, albeit with several critical modifications:
  1. 1.
    LLM-jp is built upon Llama 2 (Touvron et al., 2023b), whereas Amber is derived from LLaMA (Touvron et al., 2023a).
  2. 2.
    In OLMo-0724, LLM-jp, and Amber, layer normalization is applied before the self-attention and MLP sublayers; in OLMo-2, layer normalization is applied after these sublayers.
  3. 3.
    Regarding activation normalization, OLMo-2, LLM-jp, and Amber use RMSNorm, whereas OLMo-0724 adopts a nonparametric norm.
  4. 4.
    In all models, the output of the self-attention mechanism is added to the residual stream preceding the MLP.
  5. 5.
    In all models, the ReLU activation function is replaced by the SwiGLU activation function (Shazeer, 2020).
  6. 6.
    All models substitute absolute positional embeddings with rotary positional embeddings (Su et al., 2023).
  7. 7.
    To simplify the self-attention computations, LLM-jp uses grouped query attention (Ainslie et al., 2023).
  8. 8.
    For enhanced training stability, OLMo-2 and OLMo-0724 both use QKV Clipping.
  9. 9.
    Finally, to prevent excessively large attention logits—and consequently prevent the training loss from diverging—OLMo-2 normalizes the Key and Query projections via RMSNorm before computing the attention.
A.2 fMRI Datasets
MRI data were acquired using a 3T MAGNETOM Vida scanner (Siemens, Germany) with a standard Siemens 64-channel volume coil. Functional brain images based on the blood oxygenaction level-dependent (BOLD) signal were collected via a multiband gradient echo-planar imaging sequence (Moeller et al., 2010) (TR = 1,000 ms, TE = 30 ms, flip angle = 60, voxel size = 2 $\times$ 2 $\times$ 2 mm3, matrix size = 96 $\times$ 96, 72 slices with a thickness of 2 mm, slice gap 0 mm, FOV = 192 $\times$ 192 mm2, bandwidth 1736 Hz/pixel, partial Fourier 6/8, multiband acceleration factor 6). Anatomical data were acquired using the same 3T scanner using T1-weighted MPRAGE (TR = 2530 ms, TE = 3.26 ms, flip angle = 9, voxel size = 1 $\times$ 1 $\times$ 1 mm3, FOV = 256 $\times$ 256 mm2). The preprocessing of the fMRI data included motion correction, coregistration, and detrending. All participants provided written informed consent for this study, which was conducted under the approval of the relevant ethics and safety committee.
This dataset comprises nine videos of movies or drama series as experimental stimuli (ten episodes in total). The videos span a diverse range of genres: eight international movies or dramas and one Japanese animation. The average duration across the ten episodes is 49.98 min (minimum 21 min, maximum 125 min). Each episode is segmented into 2-9 parts, each lasting approximately 10 min. These segments were administered as fMRI stimuli.
This dataset provides three types of natural language annotations describing the stimulus videos: Objective Information, Speech Transcription, and Narrative Content (Story). Each type of annotation captures distinct semantic content relevant to narrative comprehension. We used the Narrative Content (Story) annotation for the main analysis and the Objective Information annotation for the control analysis. All annotations were originally described in Japanese. They were translated into English and back-translated into Japanese using DeepL.
A.3 Brain Encoding Models
The dataset used in this study comprises nine movies or dramas, and therefore the regularization parameters were tuned during training, using sessions from two or three movies or dramas as validation data and the remaining sessions as training data. This procedure was iterated for cross-validation. For the evaluation, we computed the Pearson’s correlation coefficients between the predicted and measured fMRI signals. Statistical significance was assessed using a blockwise permutation test. Specifically, to generate a null distribution, we shuffled the voxel’s measured response time course before calculating the Pearson’s correlation between the predicted response time course and the permuted response time course. During this process, we shuffled the measured response time course in blocks of 10 TRs to preserve the temporal correlation between slices. We identified voxels having scores significantly higher than those expected by chance in the null distribution.
On the basis of the encoding analysis results in Section 3.2, we performed an analysis using the regions of interest (ROIs) included in the DeusTex atlas. We selected focal ROIs that (1) exhibited a trend of three-phase transitions in encoding accuracy throughout training (though the specific timing of these shifts varied by participant), (2) contained voxels showing a pronounced manifestation of these three-phase transitions, and (3) demonstrated comparatively high encoding accuracy for every participant.
All encoding (Section 3.2) and probing (Section 3.3) analyses were conducted using the himalaya library [^3] https://github.com/gallantlab/himalaya (Dupré la Tour et al., 2022) and the drama2brain library [^4] https://github.com/yu-takagi/drama2brain (Nakagi et al., 2024). To extract latent representations from the MLP layers of OLMo-2, we modified the code from the Transformers library [^5] https://github.com/huggingface/transformers (Wolf et al., 2020). To extract latent representations from the MLP layers of OLMo-0724, LLM-jp, and Amber, we modified the code from the TransformerLens library [^6] https://github.com/TransformerLensOrg/TransformerLens (Nanda & Bloom, 2022). We will make our source code and training data for the encoding, probing (See Section 3.3), and benchmark (See Section 3.4) analyses publicly available on acceptance.
A.4 Downstream Datasets
For MMLU, we use the original English dataset from Hendrycks et al. (2020) and its Japanese translation from OpenAI (2024). For control analysis, we use the Chinese translation from OpenAI (2024). Each of these datasets (English/Japanese/Chinese) comprises 13,571 samples. For CSQA, we use the original English dataset from Talmor et al. (2019) and the Japanese dataset from Kurihara et al. (2022), which contain 10,957 and 8,934 samples, respectively. For ARC (both the ARC-Challenge and ARC-Easy subsets), we use the original English dataset from Clark et al. (2018) and its Japanese translation, resulting in 7,778 samples for both the English and Japanese versions. For HellaSwag, we use the original English dataset from Zellers et al. (2019) and its Japanese translation, resulting in 9,658 samples for both the English and Japanese versions. We use the OpenAI API (GPT 4o-mini) for translation.
A.5 Intrinsic Dimensions
We used GRIDE (Denti et al., 2022) to compute the IDs. GRIDE extends the TwoNN estimator (Facco et al., 2017) to general scales.
Estimation Procedure using GRIDE
GRIDE employs the following ratio as its fundamental component:
$\mu_{i,2k,k}=\frac{r_{i,2k}}{r_{i,k}}$
where $r_{i,j}$ denotes the Euclidean distance between point $i$ and its $j$-th nearest neighbor. Under the assumption of a locally uniform density distribution, these ratios $\mu_{i,2k,k}$ are shown to follow a generalized Pareto distribution:
$f_{\mu_{i,2k,k}}(\mu)=\frac{d\left(\mu^{\,d-1}\right)^{k-1}}{B(k,k)\mu^{\,d(2k-1)+1}}$
where $B(\cdot,\cdot)$ is the beta function. Furthermore, assuming independence among the ratios $\mu_{i,2k,k}$ from different points, The likelihood of this distribution can be numerically maximized to obtain the ID. In this study, for each model checkpoint, we selected the value of $k$ at which the mean ID across layers stabilized at its maximum, thereby determining the estimated IDs.
A.6 Determining the Layers of Interest
In interpreting the learning dynamics of LLMs from three distinct perspectives, we determined which layers merited attention based on (1) each layer’s encoding accuracy, (2) each layer’s probing accuracy, and (3) each layer’s benchmark accuracy. We obtained the encoding accuracy and the probing accuracy according to the methods described in Sections 3.2 and 3.3, respectively. We computed each layer’s benchmark accuracy using Logit Lens (Nostalgebraist, 2020).
Logit Lens
In the output layer of an LLM, an unembedding matrix is employed to convert vectors into tokens by projecting the hidden-layer vectors within the model onto the vocabulary dimension. A softmax function (or similar) is then applied to compute probabilities and generate the output tokens. This process is referred to as “unembedding”.
The hidden-layer vectors within the model have the same dimensionality as the vectors in the output layer, and therefore the unembedding procedure can be applied to the hidden-layer vectors, thereby gaining insight into the intermediate processes. Logit Lens is a tool specifically devised for this purpose.
Measuring Benchmark Accuracy by Layer
Using Logit Lens, we extracted the probability distribution over the final token predicted from each intermediate layer, and designated the token assigned the highest probability as that layer’s output. We then computed the proportion of correctly answered questions by dividing the number of correct answers by the total number of questions, analogous to the procedure described in Section 3.4. This proportion was treated as the benchmark accuracy for that layer.
In Section B.5, we presented the layer-wise accuracy of OLMo-2, OLMo-0724, and LLM-jp with respect to the three metrics at each training checkpoint, thereby determining which layers were to be examined in greater detail.
A.7 Examples of 5-Shot Prompts
Figures A.1, A.2, A.3, and A.4 show examples of the English and Japanese 5-shot prompts used for each downstream task described in Section 3.3.
Figure A.1: Example of MMLU (English and Japanese) prompt.
Figure A.2: Example of CSQA (English and Japanese) prompt.
Figure A.3: Example of ARC (English and Japanese) prompt.
Figure A.4: Example of HellaSwag (English and Japanese) prompt.
Appendix B Additional Results
B.1 Learning Dynamics of LLMs
Here, we present supplementary results corresponding to Section 4.1. Figures B.1, B.2, B.3, B.4, B.5, and B.6 illustrate outcomes for all participants and main LLMs (OLMo-2, OLMo-0724, LLM-jp) when using layers adjacent to those of Figure 2. We demonstrate that comparable encoding results can be obtained using these layers. Figure B.7 displays the outcomes of DM06 and the main LLMs (OLMo-2, OLMo-0724, LLM-jp) on different downstream tasks (CSQA, ARC, HellaSwag) to those of Figure 2, showing that similar benchmark and probing results are achieved for these alternate tasks. Figure B.8 presents the results obtained using Amber, a different LLM from those employed for Figure 2. We confirm a phase transition in encoding accuracy around layer 22; however, downstream task performance by Amber only extends to instruction-following, and thus the increase in probing accuracy observed in Phase 3 is slightly attenuated. Figure B.9 reports the results obtained when the LLMs are given input in Chinese, a different language to that used for Figure 2. This figure indicates that the phase transition is not observed in a language on which the model has not been trained. Finally, Figure B.10 presents the results of using the Object annotation, which differs from that used for Figure 2; here, too, we confirm a similar phase transition.
B.2 Changes in the Relationship with the Brain
In this section, we present supplementary findings related to Section 4.2. Figures B.11, B.12, B.13, B.14, B.15, and B.16 illustrate the outcomes for all participants and main LLMs (OLMo-2, OLMo-0724, LLM-jp). We show that similar results are observed for every participant. Furthermore, when the language in which the LLM was trained is used as input, the results are similar to those depicted in Figure 3 across the three LLMs.
The line graphs presented in Figures B.11B.16 illustrate how the voxel-wise changes aggregate within three major ROIs. Across participants, Phase 1 shows that there is a sharp increase in accuracy for all ROIs, while Phases 2 and 3 exhibit more localized or modest changes. Although some voxels within the temporal and occipital cortex exhibit higher prediction accuracy after Phase 3 than after Phase 1, this trend is less evident at the ROI level. This discrepancy suggests that finer-grained voxel-level analyses may be more sensitive than coarse ROI-level approaches in capturing such effects.
B.3 Evolution of LLM Internal Representations for Downstream Tasks
In this section, we present supplementary findings related to Section 4.2. Figures B.17, B.18, and B.19 correspond to Figure 4a for the main LLMs (OLMo-2, OLMo-0724, LLM-jp) and all downstream tasks (MMLU, CSQA, ARC, HellaSwag). The results confirm that when the learned language is fed into the LLM, its neurons progressively acquired robust representations pertinent to each task. Figure B.20 corresponds to Figure 4b for OLMo-2/OLMo-0724 and all downstream tasks. We observe a similar tendency in both LLMs. Additionally, we examine the relationships between HellaSwag and the other three tasks, observing neurons that specialize in both tasks as well as neurons dedicated to a single task.
Summarizing these findings alongside our primary results (Section 4.2), the neuron-wise probing accuracy for MMLU and ARC exhibits a remarkably high correlation, followed by a moderately positive correlation between HellaSwag and those two tasks (MMLU and ARC). By contrast, CSQA displays no correlation with any of the tasks. These observations suggest that the way each neuron in an LLM acquires its representations varies according to the nature of the task (e.g., required capabilities and answer formats).
B.4 Changes in the Nature of Activations
Figure B.21 shows additional results corresponding to those in Figure 5 produced by the other LLMs (OLMo-0724, LLM-jp) when provided with the learned language data. When the number of training tokens exceeds 109, the changes in IDs throughout the training processes of the other LLMs exhibit a high correlation with variations in encoding accuracy, although some exhibited distinct patterns. By contrast, when the number of training tokens is less than 109 (which is not possible for OLMo-2 or OLMo-0724), there is a precipitous drop from initially very high ID values in the activations of LLM-jp.
B.5 Layers of Interest
Figures B.22, B.23, and B.24 show the layer-wise encoding, probing, and benchmark accuracies of OLMo-2, OLMo-0724, and LLM-jp at each training checkpoint, thereby determining which layers to examine in greater detail.
We can observe the phase transition in encoding accuracy (particularly the transition from Phase 1 to Phase 2) in layers 20–28 of OLMo-2, layers 30–32 of OLMo-0724, and layers 15–28 of LLM-jp. We can further identify the phase transition in probing accuracy (specifically the transition from Phase 2 to Phase 3) in layers 19–32 of OLMo-2, layers 25–32 of OLMo-0724, and layers 19–32 of LLM-jp. Finally, we can detect the phase transition in benchmark accuracy in layers 22–32 of OLMo-2, layers 28–32 of OLMo-0724, and layers 20–32 of LLM-jp. In this study, we focused on layers in which all three of these transitions emerge. Consequently, we confirmed this tendency in the later layers, namely in layers 22–28 for OLMo-2, layers 30–32 for OLMo-0724, and layers 20–28 for LLM-jp. Moreover, because OLMo-2 and LLM-jp exhibited the three transitions most prominently at layer 25, and the transitions were most prominent for OLMo-0724 at layer 30, we present the principal analytical results for these layers.
Figure B.1: Results for all participants regarding learning dynamics of layers 24, 25, 26 of OLMo-2 exhibiting three phase transitions when using English annotation and MMLU.
Figure B.2: Results for all participants regarding learning dynamics of layers 29, 30, 31 of OLMo-0724 exhibiting three phase transitions when using English annotation and MMLU.
Figure B.3: Results for all participants regarding learning dynamics of layers 24, 25, 26 of LLM-jp exhibiting three phase transitions when using English annotation and MMLU.
Figure B.4: Results for all participants regarding learning dynamics of layers 24, 25, 26 of OLMo-2 when using Japanese annotation and MMLU.
Figure B.5: Results for all participants regarding learning dynamics of layers 29, 30, 31 of OLMo-0724 when using Japanese annotation and MMLU.
Figure B.6: Results for all participants regarding learning dynamics of layers 24, 25, 26 of LLM-jp exhibiting three phase transitions when using Japanese annotation and MMLU.
Figure B.7: Results from a single participant (DM06) for learning dynamics of layer 25 of OLMo-2 and LLM-jp, layer 30 of OLMo-0724 exhibiting three phase transitions when using other tasks.
Figure B.8: Results for all participants regarding learning dynamics of layer 22 of Amber exhibiting three phase transitions when using English/Japanese annotation and MMLU.
Figure B.9: Results for all participants regarding learning dynamics of layer 25 of OLMo-2 when using Chinese annotation and MMLU.
Figure B.10: Results for all participants regarding learning dynamics of layer 25 of OLMo-2 when using English Object annotation and MMLU.
Figure B.11: Results for all participants regarding changes in the relationship with the brain using layer 25 of OLMo-2 and English annotation and MMLU. The line graph below shows the results of the ROI-level analysis. The vertical axis denotes the average encoding accuracy of significant voxels within each ROI. The solid line represents the mean across participants, and the dashed lines represent individual participant results. Black contours show several representative, anatomically defined ROIs.
Figure B.12: Results for all participants regarding changes in the relationship with the brain using layer 30 of OLMo-0724 and English annotation and MMLU.
Figure B.13: Results for all participants regarding changes in the relationship with the brain using layer 25 of LLM-jp and English annotation and MMLU.
Figure B.14: Results for all participants regarding changes in the relationship with the brain using layer 25 of OLMo-2 and Japanese annotation and MMLU.
Figure B.15: Results for all participants regarding changes in the relationship with the brain using layer 30 of OLMo-0724 and Japanese annotation and MMLU.
Figure B.16: Results for all participants regarding changes in the relationship with the brain using layer 25 of LLM-jp and Japanese annotation and MMLU.
Figure B.17: Results for all downstream tasks regarding changes in the activations of OLMo-2.
Figure B.18: Results for all downstream tasks regarding changes in the activations of OLMo-0724.
Figure B.19: Results for all downstream tasks regarding changes in the activations of LLM-jp.
Figure B.20: Relationship between probing accuracies in OLMo-2 (layer 25) and OLMo-0724 (layer 30) across English MMLU, CSQA, ARC, and HellaSwag.
Figure B.21: Variations in correlation coefficients (left), encoding accuracy, and IDs (right) of the activations of OLMo-0724 (layer 30)/LLM-jp (layer 25) using learned languages across checkpoints.
Figure B.22: Layers of interest for OLMo-2.
Figure B.23: Layers of interest for OLMo-0724.
Figure B.24: Layers of interest for LLM-jp.

### v2

@@
 a Brain encoding analysis.
-b Probing analysis (top) and the benchmark analysis (bottom).
+b Probing analysis (top) and benchmark analysis (bottom).
 c Three phase-transition phenomena during the learning process of LLMs, as identified through the results of the encoding, probing, and benchmark analyses.
-Red, green, and blue lines indicate encoding, probing, and benchmark accuracies, respectively, accros all LLM checkpoints.
+Red, green, and blue lines indicate encoding, probing, and benchmark accuracies, respectively, across all LLM checkpoints.
 2 Related work
@@
 For this, we use four pre-trained models with available training checkpoints: OLMo-2 (OLMo et al., 2024), OLMo-0724 (Groeneveld et al., 2024), and LLM-jp (LLM-jp et al., 2024) for the main analysis, and Amber (Liu et al., 2023) for the control analysis.
+The number of checkpoints used for each model ranges from 18 to 28.
 Table 1 presents an overview of these LLMs.
 OLMo-2, OLMo-0724, and Amber are English-centric LLMs trained on publicly accessible datasets, while LLM-jp is a bilingual Japanese-English model trained on a roughly equal mix of Japanese and English.
-Each model uses a different training corpus and tokenizer, with vocabulary sizes of 100,352 (OLMo-2), 50,304 (OLMo-0724), 99,584 (LLM-jp), and 32,000 (Amber).
+Each model uses a different training corpus and tokenizer, with vocabulary sizes ranging from 32,000 to 100,352.
 These LLMs have between 6.74–7.3B parameters, 32 hidden layers, and a hidden dimension of 4,096.
 They are all based on a decoder-only Transformer architecture (Vaswani et al., 2017), but each LLM incorporates a few critical modifications.
-There are also notable differences across these LLMs in terms of their layer normalization and attention mechanisms (see Section A.1 for architectural details).
+There are also notable differences across these LLMs in terms of their layer normalization and attention mechanisms.
+See Section A.1 for further details on the models used.
 The multilayer perceptron (MLP) layers require most parameters and represent essential features (Bereska & Gavves, 2024; Geva et al., 2021).
 Thus, we use their neural activation as the internal representations of each model.
-We use 28 checkpoints for OLMo-2 (1B-3,896B training tokens), 23 for OLMo-0724 (4B–2,724B), 27 for LLM-jp (4.2M–1,258B), and 18 for Amber (3.5B–1,259B).
-See Section A.1 for further details on the model sources and specific checkpoints.
 3.2 Brain encoding models
 Our initial approach for analyzing the learning dynamics of the LLMs involves an investigation of how their activations progressively align with brain activity during the training process.
-Specifically, for each checkpoint described in Section 3.1, we perform an encoding analysis by evaluating the prediction accuracy of a learned linear mapping from each layer’s activations to brain activity (Naselaris et al., 2011; Nishimoto et al., 2011; Huth et al., 2012) (Figure 1a).
-These analysis are conducted separately for each participant.
+Specifically, for each checkpoint described in Sections 3.1 and A.1, we perform an encoding analysis by evaluating the prediction accuracy of a learned linear mapping from each layer’s activations to brain activity (Naselaris et al., 2011; Nishimoto et al., 2011; Huth et al., 2012) (Figure 1a).
+These analyses are conducted separately for each participant.
 fMRI datasets
@@
 This dataset provides brain activity data from six healthy participants with normal or corrected-normal vision (three females; ages 22–40, mean = 28.7), while they freely watched 8.3 hours of movies or drama series inside a 3T functional Magnetic Resonance Imaging (fMRI) scanner.
-All participants were right-handed and native Japanese speakers.
+All participants were native Japanese speakers.
 The dataset includes nine video clips of movies or drama series as stimuli: eight international titles and one Japanese title.
@@
 We use the Narrative Content annotation, which describes the background story of the scene at 5-s intervals, for the main analysis and the Objective Information annotation, which describes the objects in the scene every second, for the control analysis, both in English and Japanese.
-Section A.2 provides additional details regarding the dataset.
+See Section A.2 for additional dataset details.
 In this study, we use 29,993 seconds of data covering all six participants.
@@
 Specifically, we use the fMRI scanning sessions corresponding to the last split of each movie or drama series (7,737 seconds in total) as the test data.
-The remainder of the sessions (22,262 seconds in total) constitutes the training data.
+The remainder of the sessions (22,262 seconds in total) forms the training data.
 Model construction
@@
 In the main analysis, we focus on later layers within each LLM.
-Specifically, we focus on layer 25 in both OLMo-2 and LLM-jp, and layer 30 in OLMo-0724.
-These layers show the most substantial checkpoint-wise changes and are strongly involved in the emergent behavior of each model, as verified in all three evaluations (encoding, probing, and benchmark analyses; see Section A.6 for details).
+Specifically, we focus on layer 25 in both OLMo-2 and LLM-jp, and layer 30 in OLMo-0724, which show the greatest checkpoint-wise changes and are strongly involved in each model’s emergent behavior across all evaluations (encoding, probing, and benchmark analyses; see Section A.6 for details).
 We confirm that the layers surrounding the main focal layer exhibit similar patterns of results (see Figure B.5).
@@
 To this end, we perform a probing analysis by evaluating how well linear models can predict neuron activations from the answer labels of downstream tasks across all hidden layers (Figure 1b, top).
-These analysis are conducted separately for each LLM checkpoint.
+These analyses are conducted separately for each LLM checkpoint.
 Downstream datasets
 We use four downstream tasks: Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2020), CommonsenseQA (CSQA) (Talmor et al., 2019), AI2 Reasoning Challenge (ARC) (Clark et al., 2018), and HellaSwag (Zellers et al., 2019).
-MMLU assesses broad knowledge and problem-solving abilities using multidisciplinary coverage of 57 subjects, CSQA tests everyday conceptual commonsense reasoning, ARC probes elementary-level scientific knowledge, and HellaSwag assesses contextual commonsense reasoning in typical scenarios.
 For both the probing and benchmark analyses (see Section 3.4), we use 5-shot prompts in both English and Japanese.
 We exclude prompts exceeding each model’s maximum context length.
-See Section A.4 for additional details about downstream datasets.
+See Section A.4 for additional details about downstream datasets and Section A.7 for examples of the prompts used in these datasets.
 In the probing analysis, each dataset is split into training and test datasets at a 4:1 ratio.
-Because MMLU comprises multiple subject areas, we split the dataset by subject.
 All prediction performance results are computed using the test dataset.
-Examples of the prompts used in these datasets are presented in Section A.7.
 Probing for MLP activations
@@
 For each sample in the downstream tasks, we create an answer matrix (the number of samples $\times$ the number of choices), where the correct and incorrect choices are labeled as 1 and 0, respectively.
-Our objective is to learn a mapping from this answer matrix to the neuron activations across all layers of the LLMs.
-To this end, we use L2-regularized linear regression to estimate the linear weights that transform the answer matrix corresponding to the training data into the observed neuron activations in the LLMs, and then apply these learned weights to the test data.
+Our objective is to learn a mapping from this answer matrix to the activations across all layers of the LLMs.
+To this end, we use L2-regularized linear regression to estimate the linear weights that transform the answer matrix corresponding to the training data into the observed activations in the LLMs, and then apply these learned weights to the test data.
 The regularization parameter is optimized via 4-fold cross-validation for each neuron in the training dataset.
-To mitigate the subject-based bias of MMLU, we shuffle the training indices, and then perform cross-validation to ensure balanced distributions in each fold.
 We then evaluate the prediction accuracy by computing the Pearson’s correlation coefficients between the predicted and actual activations.
@@
 4 Results
+4.1 Learning dynamics of LLMs
 Figure 2: Learning dynamics of LLMs exhibiting three phase transitions.
@@
 No checkpoints preceding the 109 training tokens have been made publicly available aside from LLM-jp.
-4.1 Learning Dynamics of LLMs
 […]
@@
 Brain regions with significant accuracy at either of the two checkpoints are colored ($p<0.05$, FDR-corrected).
-Voxels that exhibit higher accuracy at the later checkpoint are indicated in red, whereas those that exhibit higher accuracy at the earlier checkpoint are indicated in blue.
+Voxels exhibiting higher accuracy at the later checkpoint are indicated in red, whereas those exhibiting higher accuracy at the earlier checkpoint are indicated in blue.
 4.2 What happens during each phase?
@@
 Notably, these patterns only emerge when the input language matches one on which the model has been trained.
-Further details for all participants, ROIs, other phases, and LLMs (OLMo-2, OLMo-0724, LLM-jp) can be found in Figures B.11B.16.
+Further details for all participants, ROIs, other phases, and LLMs can be found in Figures B.11B.16.
 Figure 4: Probing results.
@@
 The legends corresponds to the number of training tokens.
-b Relationship between probing accuracy in the neurons of OLMo-2 across English MMLU, CSQA, and ARC.
+b Relationship between probing accuracy in the neurons of OLMo-2 (layer 25) across English MMLU, CSQA, and ARC.
 Each axis denotes the probing accuracy for the respective task, and the color gradient reflects the number of training tokens.
@@
 We now investigate how the neuron-level internal representations evolve over the course of training.
-For this purpose, we examine changes in the distribution of probing accuracy for each neuron at the early, middle, and final checkpoints (105B, 1259B, 3896B training tokens in OLMo-2) of Phase 3.
+For this purpose, we examine changes in the distribution of probing accuracy for each neuron at the early, middle, and final checkpoints (302B, 1259B, 3896B training tokens in OLMo-2) of Phase 3.
 Figure 4a presents the results for neurons in layers 5, 15, 25, and 30 of OLMo-2 when using MMLU.
@@
 Figures B.17B.19 show the results corresponding to all downstream tasks for OLMo-2, OLMo-0724, and LLM-jp, confirming that a similar trend emerges across every downstream task when each model is provided with the language on which it has been trained.
-Figure 5: The Nature of Activations.
-a Variations in correlation coefficients of the activations of OLMo-2 across checkpoints.
-b IDs (purple line) and average encoding accuracy for all voxels of a single participant (DM06) (red line) across checkpoints.
 We can also explore whether different sets of neurons acquire specialized representations for distinct downstream tasks, focusing on the same checkpoints examined in Figure 4a.
@@
 We now examine whether these alterations reflect fundamental shifts in the activations of the LLMs themselves.
+Figure 5: The Nature of Activations.
+a Variations in correlation coefficients of the activations of OLMo-2 across checkpoints.
+b IDs (purple line) and average encoding accuracy for all voxels of a single participant (DM06) (red line) across checkpoints.
 As shown in Figure 5a, within-phase activations are highly similar, but substantial changes emerge at each phase transition.
@@
 Next, we quantify the dimensionality of these activations using the IDs defined by GRIDE.
-Figure 5b shows that there is a strong similarity between the brain encoding accuracy and IDs during Phases 1 and 2, consistent with earlier findings that single-checkpoint encoding accuracy and dimensionality are highly correlated (Cheng & Antonello, 2024).
+Figure 5b shows that there is a strong similarity between the brain encoding accuracy and IDs during Phases 1 and 2, consistent with earlier findings that single-checkpoint encoding accuracy and IDs are highly correlated (Cheng & Antonello, 2024).
 Our findings extend this result by demonstrating a similarly robust correlation across multiple checkpoints.
-Nevertheless, the brain encoding accuracy and dimensionality are not perfectly congruent (see Phase 3), leaving open questions about the aspects of activation changes that intrinsic dimensionality alone does not capture.
+Nevertheless, the brain encoding accuracy and IDs are not perfectly congruent (see Phase 3), leaving open questions about the aspects of activation changes that intrinsic dimensionality alone does not capture.
 5 Discussion and conclusions
-In this study, we interpreted the learning dynamics of LLMs from three perspectives: alignment between the brain and the LLMs, internal representations associated with downstream tasks, and actual downstream task performance.
+In this study, we interpreted the learning dynamics of LLMs from three perspectives: alignment between LLMs and the brain, internal states associated with downstream tasks, and downstream task performance.
 Our novel approach elucidated a three-stage phase transition—including the precise timing of each transition—that emerges consistently during training, even among LLMs that diverge substantially in both training data and architecture.
@@
 Additionally, the learning dynamics that we have revealed are reminiscent of the “synaptic overproduction and pruning” posited in neuroscience, where an initial synaptic overgrowth eventually refines the neural circuits (Peter R., 1979; Bourgeois et al., 1994).
-These transient changes in the brain suggest the potential for more efficient neural information processing, leading to enchanced cognitive function.
+These transient changes in the brain suggest the potential for more efficient neural information processing, leading to enhanced cognitive function.
 Thus, a promising avenue for future work involves examining whether the phenomenon observed in this study arises from the LLM pruning superfluous internal representations in pursuit of more refined ones.
@@
 Moreover, alignment with semantic regions in the temporal cortex generally rises over the course of training, implying a dynamic reorganization in how models correspond to brain activity.
-By capturing shifts in the activations of LLM neurons pertaining to downstream tasks, we elucidated how task-specific neurons form and how much specialization they share among tasks.
+By capturing shifts in LLM neuron activations related to downstream tasks, we elucidated how task-specific neurons form and how much specialization they share among tasks.
 These findings align with previous studies showing neurons that encode particular linguistic information (Tenney et al., 2019; Dai et al., 2022; Wang et al., 2022; Gurnee et al., 2023).
 We observed that neurons contributing substantially to certain tasks emerge abruptly in later layers as training progresses.
-Furthermore, the variations in neuron sharing across tasks suggest that differences in activation patterns—reflecting task-specific characteristics (e.g., required capabilities, difficulty levels, and answer formats)—are indeed discernible at the neuronal level.
+Moreover, the variations in neuron sharing across tasks suggest that differences in activation patterns—reflecting task-specific characteristics (e.g., required capabilities, difficulty levels, and answer formats)—are indeed discernible at the neuronal level.
 Taken together, our results offer a more comprehensive view of LLM neurons during the learning process, building on past work that focused solely on trained models.
@@
 Tables 1 and 2 present an overview of the LLMs and detailed information on their respective training checkpoints used in this study.
+We used 28 checkpoints for OLMo-2 (1B-3,896B training tokens), 23 for OLMo-0724 (4B–2,724B), 27 for LLM-jp (4.2M–1,258B), and 18 for Amber (3.5B–1,259B).
 In selecting these checkpoints, we took particular care to ensure that the number of training tokens was as closely aligned as possible across the four LLMs.
-Model Layers Width Params. Trn. Tokens (Ckpts.)
-OLMo-2 32 4096 7.3B 1B - 3896B (28)
-OLMo-0724 32 4096 6.89B 4B - 2724B (23)
-LLM-jp 32 4096 7.29B 4.2M - 1258B (27)
-Amber 32 4096 6.74B 3.5B - 1259B (18)
+Model Layers Width Params. Vocab. sizes Trn. Tokens (Ckpts.)
+OLMo-2 32 4096 7.3B 100352 1B - 3896B (28)
+OLMo-0724 32 4096 6.89B 50304 4B - 2724B (23)
+LLM-jp 32 4096 7.29B 99584 4.2M - 1258B (27)
+Amber 32 4096 6.74B 32000 3.5B - 1259B (18)
 Table 1: Overview of LLMs
@@
 MRI data were acquired using a 3T MAGNETOM Vida scanner (Siemens, Germany) with a standard Siemens 64-channel volume coil.
-Functional brain images based on the blood oxygenaction level-dependent (BOLD) signal were collected via a multiband gradient echo-planar imaging sequence (Moeller et al., 2010) (TR = 1,000 ms, TE = 30 ms, flip angle = 60, voxel size = 2 $\times$ 2 $\times$ 2 mm3, matrix size = 96 $\times$ 96, 72 slices with a thickness of 2 mm, slice gap 0 mm, FOV = 192 $\times$ 192 mm2, bandwidth 1736 Hz/pixel, partial Fourier 6/8, multiband acceleration factor 6).
+Functional brain images based on the blood oxygenation level-dependent (BOLD) signal were collected via a multiband gradient echo-planar imaging sequence (Moeller et al., 2010) (TR = 1,000 ms, TE = 30 ms, flip angle = 60, voxel size = 2 $\times$ 2 $\times$ 2 mm3, matrix size = 96 $\times$ 96, 72 slices with a thickness of 2 mm, slice gap 0 mm, FOV = 192 $\times$ 192 mm2, bandwidth 1736 Hz/pixel, partial Fourier 6/8, multiband acceleration factor 6).
 Anatomical data were acquired using the same 3T scanner using T1-weighted MPRAGE (TR = 2530 ms, TE = 3.26 ms, flip angle = 9, voxel size = 1 $\times$ 1 $\times$ 1 mm3, FOV = 256 $\times$ 256 mm2).
 The preprocessing of the fMRI data included motion correction, coregistration, and detrending.
-All participants provided written informed consent for this study, which was conducted under the approval of the relevant ethics and safety committee.
+All participants are right-handed, native Japanese speakers and provided written informed consent for this study, which was conducted under the approval of the relevant ethics and safety committee.
 This dataset comprises nine videos of movies or drama series as experimental stimuli (ten episodes in total).
@@
 A.4 Downstream datasets
+MMLU assesses broad knowledge and problem-solving abilities using multidisciplinary coverage of 57 subjects, CSQA tests everyday conceptual commonsense reasoning, ARC probes elementary-level scientific knowledge, and HellaSwag assesses contextual commonsense reasoning in typical scenarios.
 For MMLU, we use the original English dataset from Hendrycks et al.
@@
 We use the OpenAI API (GPT 4o-mini) for translation.
+In the probing analysis, each dataset is split into training and test datasets at a 4:1 ratio.
+Because MMLU comprises multiple subject areas, we split the dataset by subject.
+Furthermore, during the optimization of regularization parameters described in Section 3.3, to mitigate the subject-based bias of MMLU, we shuffle the training indices, and then perform cross-validation to ensure balanced distributions in each fold.
 A.5 Intrinsic dimensions